A man invests a portion of $3,600 at 8% annual interest and the balance at 7%. If he receives $264 interest during the first year, how much did he invest at each rate?
Answer by checkley77(12569)

.08X+.07(3600-X)=264
.08X+252-.07X=264
.01X=264-252
.01X=12
X=12/.01
X=1200 INVESTED @ 8%
3600-1200=2400 INVESTED @ 7%
PROOF:
.08*1200+.07*2400=264
96+168=264
264=264